# Multi-Draft Speculative Sampling: Canonical Architectures and Theoretical Limits

### v1

@@ Appendix A Proof of Theorem 1
We now consider the problem of optimizing the acceptance probability for any given $p(\cdot)$ and $q(\cdot)$ in the general setting. Following the framework in Sun et al. [2024b], we seek to find $p_{Y|X_{1},\ldots X_{K}}(y|x_{1},\ldots,x_{K})$ for each $y,x_{1},\ldots,x_{K}\in\Omega$ such that we maximize
@@
$=\sum_{x_{1}^{K}\in\Omega_{K}}\sum_{y\in\Omega}p_{Y|X_{1}^{K}}(y|x_{1}^{K}){\mathbb{I}}_{y}(x_{1},\ldots,x_{K})\left(\prod_{i=1}^{K}p(x_{i})\right)$ (28)
$=\sum_{y\in\Omega}\sum_{x_{1}^{K}\in\Omega_{K}}p_{Y|X_{1}^{K}}(y|x_{1}^{K}){\mathbb{I}}_{y}(x_{1},\ldots,x_{K})\left(\prod_{i=1}^{K}p(x_{i})\right)$ (29)
@@
A.1 Extension to Non-IID Setting
The proof in Theorem 1 assumed that $x_{1},\ldots,X_{K}$ are sampled form the same underlying distribution $p(\cdot)$. Here we provie a natural extension when the sampled are still independently sampled from from a non-i.i.d. distribution i.e $X_{i}\sim p_{i}(\cdot)$.
@@ A.1 Extension to Non-IID Setting
Let $P^{\star}(\mathrm{acc})$ be the acceptance probability for the optimal token level selection rule when ${\mathcal{S}}\sim\prod_{i=1}^{K}p_{i}(X_{i})$. Then we have
$P^{\star}(\mathrm{acc})=\max_{\{\beta_{y}(x_{1}^{K})\}}\left\{\sum_{y\in\Omega}\min\left(q(y),\sum_{x_{1},\ldots,x_{K}\in\Omega}\beta_{y}(x_{1}^{K})\cdot\prod_{i=1}^{K}p_{i}(x_{i})\right)\right\}$ (38)
where the maximum is over $\beta_{y}(x_{1}^{K})$ for each $\{x_{1},\ldots,x_{K},y\}\in\Omega$ such that $0\leq\beta_{y}(x_{1}^{K})\leq 1$, and
@@
$\beta_{y}(x_{1}^{K})=0,\quad y\notin\{x_{1},\ldots,x_{K}\}.$ (40)
Furthermore if $\{\beta^{\star}_{y}(x_{1}^{K})\}$ denotes the parameters that achieve the maximum in (38), then $P^{\star}(\mathrm{acc})$ can be attained by a two step approach as follows: in the first step, given the list of input tokens $\{x_{1},\ldots,x_{K}\}$, we apply Importance Weighted Speculative Sampling in Definition 1 with parameters $\beta^{\star}_{y}(x_{1},\ldots,x_{K})$ to output an intermediate token $y\in\{x_{1},\ldots,x_{K}\}$; in the second step we apply a single-draft speculative sampling scheme [Chen et al., 2023, Leviathan et al., 2023] on the selected token $y$ to generate the final output token.
The proof of Theorem 4 is identical to the proof of Theorem 1. We note that replacing the distribution of ${\mathcal{S}}$ from $\prod_{i=1}^{K}p(X_{i})$ to $\prod_{i=1}^{K}p_{i}(X_{i})$ does not affect any of the steps in the original proof.
Appendix B Proof of Theorem 2
@@
Our initial attempt was to directly apply Fourer-Motzkin (FM) elimination technique Ziegler [2012], Dantzig and Curtis Eaves [1973] to (41). However a direct application of FM elimination does not appear to be tractable for arbitrary sized alphabets, as the elimination of each variable introduces a large number of inequalities. Our key observation is that (41) is equivalent to the following relaxed set of inequalities:
@@
Appendix C Connection between Theorem 2 and Polyhedral Cone Representation
@@
The Minikowski-Weyl Theorem [Fukuda and Prodon, 1995] guarantees that for every ${\mathbf{B}}$ in (106) there exists a ${\mathbf{R}}$ in (107) of finite dimensions such that ${\mathcal{B}}={\mathcal{R}}$. Furthermore the double-description method is an algorithmic way of computing ${\mathbf{R}}$ given ${\mathbf{B}}$ and vice versa. Using the package skeleton for double description [Zny, 2018] we could show that for the ${\mathbf{B}}$ matrix in (105) the associated $R$ matrix is given by:
@@
In the last step we use the fact that $\sum q_{i}=\sum p_{i}=1$. It thus follows that ${\mathbf{R}}^{T}{\mathbf{b}}\geq 0$ if and only if $q_{i}\geq p_{i}^{2}$ and $q_{i}+q_{j}\geq(p_{i}+p_{j})^{2}$ holds as stated in Theorem 2. Thus this approach provides an alternative proof for Theorem 2 for the case of $|\Omega|=3$. We did not however find a simple approach to analytically compute the generator representation ${\mathcal{R}}$ from the hyperplane representation ${\mathcal{B}}$ for arbitrary dimensions. On the other hand we used the numerical implementation of the double description method to compute ${\mathbf{B}}$ and ${\mathbf{R}}$ for the case of up-to $K=6$ drafts and $|\Omega|\leq 14$ and demonstrate that the natural counterpart of our result in Theorem 2 appears to be valid in all these cases
Appendix D Proof of Theorem 3
@@
Appendix E Proof of Equation (14)
@@
Appendix F Histogram of Alphabet Size for OPT model
We present additional evidence that after top-p sampling the effective alphabet size can be significantly reduced. We consider the OPT model and report the histogram of the alphabet size under the draft and target models for the XSum task in Fig. 5, the Dolly task in Fig. 6, and the WMT task in Fig. 7.
@@ Appendix F Histogram of Alphabet Size for OPT model
Appendix G LP and Fast LP version for non-identical draft distributions, $K=2$
@@
Appendix H Additional Experimental Results
H.1 Draft Acceptance Probability
In this experiment, reported in Table 3, we compare the token-level acceptance probability across different methods. We first generate a sequence of tokens via auto-regressive decoding using the target model. At each step we compute the logits generated by the draft and the target model, which we then use to compute the acceptance probability for different schemes. We average the acceptance probability over $100$ instances. We report the acceptance probabilities for the case of $K=2,4,8$ drafts.
We also consider the theoretically optimal acceptance probability for the case of $K=2$ drafts and also compute the natural extension of this expression for $K=4,8$ discussed in Remark 2 (reported using the gray font).
In the importance sampling (IS) scheme, we use a truncated linear program with a cut-off threshold $s=7$. We use top-$k$ sampling with $k=5$ for both models in this experiment. In the Dolly task we use a sampling temperature of $1.0$ for both the target and draft models. In the XSum task, we use a temperature of $0.9$ for the target and a temperature of $0.3$ for the draft model.
Note that the acceptance probability increases as we increase the number of drafts in each case. We further observe that our proposed IS scheme performs consistently better than SpecTr and SpecInfer in all cases in Table 3.
Table 3: Comparison of average acceptance probability across different tasks for $K=2,4,8$ drafts.
Scheme XSum Dolly
K=$2$ K=$4$ K=$8$ K=$2$ K=$4$ K=$8$
Optimal 0.5009 0.5226 0.5419 0.6384 0.6731 0.6962
IS 0.4933 0.5145 0.5333 0.6348 0.6691 0.6919
SpecTr 0.4889 0.5083 0.5263 0.6246 0.6560 0.6800
SpecInfer 0.4875 0.5058 0.5227 0.6202 0.6489 0.6722
H.2 Top-$k$ Sampling with Identical Draft Models
Table 4 compares the block efficiencies for different methods using $K=2$ and $K=3$ drafts. We apply top-$k$ sampling with $k=10$ and $k=5$ and use a temperature of $1.0$ for both models. In our IS scheme, we truncate the variables in the LP to $s=7$. While our proposed scheme still achieves a higher block efficiency over prior works, the gains are relatively small. This can be explained by noting that in this setting the acceptance probability of the baseline schemes is already close to the theoretical limit, as noted in Table 3.
Sampling Scheme $K=2$ drafts $K=3$ drafts
Block Efficiency Loss Block Efficiency Loss
top-$k$ ($k=10$) IS $2.48\pm 0.01$ $2.59\pm 0.02$
SpecTr $2.43\pm 0.01$ $98\%$ $2.55\pm 0.01$ $98\%$
SpecInfer $2.38\pm 0.02$ $96\%$ $2.49\pm 0.02$ $96\%$
top-$k$ ($k=5$) IS $2.52\pm 0.02$ $2.63\pm 0.03$
SpecTr $2.48\pm 0.02$ $98\%$ $2.56\pm 0.03$ $97\%$
SpecInfer $2.47\pm 0.01$ $98\%$ $2.55\pm 0.04$ $97\%$
H.3 ROUGE-L and BLEU Scores
@@
In the experiments involving the XSum and WMT tasks we measure the ROUGE-L [Lin, 2004] and BLEU [Papineni et al., 2002] scores respectively. Table 5 and Table 6 show the ROUGE-L and BLEU scores for the case of identical draft models, corresponding to the experiment in Figure 3 in Section 5.1.

### v2

@@ Appendix B Proof of Theorem 1
We now consider the problem of optimizing the acceptance probability for any given $p(\cdot)$ and $q(\cdot)$ in the general setting. Following the framework in Sun et al. (2024b), we seek to find $p_{Y|X_{1},\ldots X_{K}}(y|x_{1},\ldots,x_{K})$ for each $y,x_{1},\ldots,x_{K}\in\Omega$ such that we maximize
@@
$=\sum_{x_{1}^{K}\in\Omega^{K}}\sum_{y\in\Omega}p_{Y|X_{1}^{K}}(y|x_{1}^{K}){\mathbb{I}}_{y}(x_{1},\ldots,x_{K})\left(\prod_{i=1}^{K}p(x_{i})\right)$ (29)
$=\sum_{y\in\Omega}\sum_{x_{1}^{K}\in\Omega^{K}}p_{Y|X_{1}^{K}}(y|x_{1}^{K}){\mathbb{I}}_{y}(x_{1},\ldots,x_{K})\left(\prod_{i=1}^{K}p(x_{i})\right)$ (30)
@@
B.1 Extension beyond IID Setting
The proof in Theorem 1 assumed that $x_{1},\ldots,x_{K}$ are sampled form the same underlying distribution $p(\cdot)$. Here we provide a natural extension when the tokens are sampled are not sampled from a product distribution but instead from a joint distribution: $p(x_{1},\ldots x_{K})$.
@@ B.1 Extension beyond IID Setting
Let $P^{\star}(\mathrm{acc})$ be the acceptance probability for the optimal token level selection rule when ${\mathcal{S}}\sim p(x_{1},\ldots x_{K})$. Then we have
$P^{\star}(\mathrm{acc})=\max_{\{\beta_{y}(x_{1:K})\}}\left\{\sum_{y\in\Omega}\min\left(q(y),\sum_{x_{1},\ldots,x_{K}\in\Omega}\beta_{y}(x_{1:K})\cdot p(x_{1:K})\right)\right\}$ (39)
where the maximum is over $\beta_{y}(x_{1:K})$ such that $0\leq\beta_{y}(x_{1:K})\leq 1$, and
$\sum_{x_{1:K}\in\Omega^{K}}\beta_{y}(x_{1:K})=1,\quad\forall y\in\Omega,$ (40)
@@
$\beta_{y}(x_{1:K})=0,\quad y\notin\{x_{1},\ldots,x_{K}\}.$ (41)
Furthermore if $\{\beta^{\star}_{y}(x_{1:K})\}$ denotes the parameters that achieve the maximum in (39), then $P^{\star}(\mathrm{acc})$ can be attained by a two step approach as follows: in the first step, given the list of input tokens $\{x_{1},\ldots,x_{K}\}$, we apply Importance Weighted Speculative Sampling in Definition 1 with parameters $\beta^{\star}_{y}(x_{1},\ldots,x_{K})$ to output an intermediate token $y\in\{x_{1},\ldots,x_{K}\}$; in the second step we apply a single-draft speculative sampling scheme (Chen et al., 2023, Leviathan et al., 2023) on the selected token $y$ to generate the final output token.
The proof of Theorem 4 is identical to the proof of Theorem 1. We note that replacing the distribution of ${\mathcal{S}}$ from $\prod_{i=1}^{K}p(X_{i})$ to the joint distribution $p(x_{1},\ldots x_{K})$ does not affect any of the steps as we did not use the property that the joint distribution factorizes in to the product.
Appendix C Proof of Theorem 2
@@
Our initial attempt was to directly apply Fourer-Motzkin (FM) elimination technique Ziegler (2012), Dantzig and Curtis Eaves (1973) to (42). However a direct application of FM elimination does not appear to be tractable for arbitrary sized alphabets, as the elimination of each variable introduces a large number of inequalities. Our key observation is that (42) is equivalent to the following relaxed set of inequalities:
@@
Appendix D Connection between Theorem 2 and Polyhedral Cone Representation
@@
The Minikowski-Weyl Theorem (Fukuda and Prodon, 1995) guarantees that for every ${\mathbf{B}}$ in (107) there exists a ${\mathbf{R}}$ in (108) of finite dimensions such that ${\mathcal{B}}={\mathcal{R}}$. Furthermore the double-description method is an algorithmic way of computing ${\mathbf{R}}$ given ${\mathbf{B}}$ and vice versa. Using the package skeleton for double description (Zny, 2018) we could show that for the ${\mathbf{B}}$ matrix in (106) the associated $R$ matrix is given by:
@@
In the last step we use the fact that $\sum q_{i}=\sum p_{i}=1$. It thus follows that ${\mathbf{R}}^{T}{\mathbf{b}}\geq 0$ if and only if $q_{i}\geq p_{i}^{2}$ and $q_{i}+q_{j}\geq(p_{i}+p_{j})^{2}$ holds as stated in Theorem 2. Thus this approach provides an alternative proof for Theorem 2 for the case of $|\Omega|=3$. We did not however find a simple approach to analytically compute the generator representation ${\mathcal{R}}$ from the hyperplane representation ${\mathcal{B}}$ for arbitrary dimensions. On the other hand we used the numerical implementation of the double description method to compute ${\mathbf{B}}$ and ${\mathbf{R}}$ for the case of up-to $K=6$ drafts and $|\Omega|\leq 14$ and demonstrate that the natural counterpart of our result in Theorem 2 appears to be valid in all these cases.
D.1 Intuition for the case when $K=3$ and $|\Omega|=3$
We consider the case of $\Omega=\{1,2,3\}$ and $K=3$ drafts. We let ${\mathbf{q}}=(q_{1},q_{2},q_{3})$ and ${\mathbf{p}}=(p_{1},p_{2},p_{3})$ denote the probabilities of the target and draft models. We build upon the ideas in Example 2 in Section C and provide some intuition that the necessary and sufficient condition for the acceptance probability to equal $1$ is:
$q_{i}>p_{i}^{3},\quad q_{i}+q_{j}\geq(p_{i}+p_{j})^{3},\qquad\forall i,j\in\{1,2,3\},i\neq j.$
We let $\alpha_{1,1,2}(1)=\Pr(Y=1,\{X_{1},X_{2},X_{3}\}=\{1,1,2\})$ i.e., the probability that the selected token equals $1$ and the input tokens are $\{1,1,2\}$ in any order. Likewise we let $\alpha_{1,1,2}(2)=\Pr(Y=2,\{X_{1},X_{2},X_{3}\}=\{1,1,2\})$. Note that $\alpha_{1,1,2}(1)+\alpha_{1,1,2}(2)=3p_{1}^{2}p_{2}$ is the probability of observing the input tokens to equal $\{1,1,2\}$ in some order. We define $\alpha_{i,i,j}(i)$ and $\alpha_{i,i,j}(j)$ for other values of $i,j$ in a similar fashion. Finally we let $\alpha_{1,2,3}(i)=\Pr(Y=i,\{X_{1},X_{2},X_{3}\}=\{1,2,3\})$ for each $i\in\{1,2,3\}$ be the probability that token $i$ is selected and the input tokens are $\{1,2,3\}$ in some order. We observe that $\sum_{i}\alpha_{1,2,3}(i)=6p_{1}p_{2}p_{3}$. We note that the probability that the selected token $Y=1$ is given by:
$p_{I}(1)=p_{1}^{3}+\alpha_{1,1,2}(1)+\alpha_{1,2,2}(1)+\alpha_{1,1,3}(1)+\alpha_{1,3,3}(1)+\alpha_{1,2,3}(1)$ (111)
Here the right hand side denotes all the events that a token $1$ appears among the in put tokens and that token $1$ is selected. Following the same relaxation as in Example 2 we can show that the acceptance probability equals $1$ if
$q_{1}\geq p_{I}(1)=p_{1}^{3}+\alpha_{1,1,2}(1)+\alpha_{1,2,2}(1)+\alpha_{1,1,3}(1)+\alpha_{1,3,3}(1)+\alpha_{1,2,3}(1)$ (112)
$q_{2}\geq p_{I}(2)=p_{2}^{3}+\alpha_{1,1,2}(2)+\alpha_{1,2,2}(2)+\alpha_{2,2,3}(2)+\alpha_{2,3,3}(2)+\alpha_{1,2,3}(2)$ (113)
$q_{3}\geq p_{I}(3)=p_{3}^{3}+\alpha_{1,1,3}(3)+\alpha_{1,3,3}(3)+\alpha_{2,2,3}(3)+\alpha_{2,3,3}(3)+\alpha_{1,2,3}(3)$ (114)
Following Example 2 we also consider the following augmented system of inequalities:
$q_{1}+q_{2}\geq(p_{1}+p_{2})^{3}+\alpha_{1,1,3}(1)+\alpha_{1,3,3}(1)+\alpha_{1,2,3}(1)+\alpha_{2,2,3}(2)+\alpha_{2,3,3}(2)+\alpha_{1,2,3}(2)$ (115)
$q_{1}+q_{3}\geq(p_{1}+p_{3})^{3}+\alpha_{1,1,3}(1)+\alpha_{1,3,3}(1)+\alpha_{1,2,3}(1)+\alpha_{2,2,3}(3)+\alpha_{2,3,3}(3)+\alpha_{1,2,3}(3)$ (116)
$q_{2}+q_{3}\geq(p_{2}+p_{3})^{3}+\alpha_{1,1,2}(2)+\alpha_{1,2,2}(2)+\alpha_{1,2,3}(2)+\alpha_{1,1,3}(3)+\alpha_{1,3,3}+\alpha_{1,2,3}(3)$ (117)
We note that (115) is a direct consequence of adding (112) and (113) and using the condition that $\alpha_{i,i,j}(i)+\alpha_{i,i,j}(j)=3p_{i}^{2}p_{j}$. The inequalities (116) and (117) follow in a similar manner. Although these inequalities are redundant their presence aids in simplifying the Fourier-Motzkin elimination as in the case of $K=2$ drafts. We do not considering summing all the inequalities in (112)-(114) as this implies $q_{1}+q_{2}+q_{3}\geq(p_{1}+p_{2}+p_{3})^{3}$, which holds trivially as both sides are equal to $1$. In order to find conditions when the probability equals $1$, we need to apply Fourier-Motzkin elimination to the system of inequalities in (112)-(117), where the variables satisfy the conditions that
$\alpha_{i,i,j}(i)+\alpha_{i,i,j}(j)=3p_{i}^{2}p_{j}$ (118)
and
$\alpha_{1,2,3}(1)+\alpha_{1,2,3}(2)+\alpha_{1,2,3}(3)=6p_{1}p_{2}p_{3}.$ (119)
and all the variables are non-negative. By following elimination procedure we can show that, similar to Example 2 and Lemma 2, the elimination of each variable in this augmented system simply amounts to dropping it from the right hand side in the system of inequalities in (112)-(117). Thus the intuition behind our conjecture is that when one considers an augmented system of inequalities, only terms of the form $\left(\sum_{s\in{\mathcal{S}}}p_{s}\right)^{K}$ appear and lower order terms do not appear. After application of the Fourier-Motzkin elimination the required condition follows.
Appendix E Proof of Theorem 3
@@
Appendix F Proof of Equation (12)
@@
Appendix G Computational Complexity of Truncated LP
Algorithm 2 Truncated LP
2:  Output: Selected token $Y_{I}$, output distribution $p_{I}(\cdot)$.
4:  Set $\Omega_{1}=\{1,\ldots,s\}$ and $\Omega_{2}=\{s+1,\ldots,n\}$.
  • Constraints: $w_{i,j}\geq 0$, $w_{i,j}+w_{j,i}=1$
7:  Compute $p_{I}(\cdot)$ according to (11).
8:  if $X_{1}=X_{2}$ then
9:     Set $Y_{I}=X_{1}$.
10:  else
11:     Let $\{X_{1},X_{2}\}=\{i,j\}$ and $i<j$.
12:     Set $Y_{I}=i$ with probability $w_{i,j}$, and $Y_{I}=j$ otherwise.
13:  end if
We study the computational complexity of the truncated linear program (see Algorithm 2) and propose a variation that could improve it further. Let $\Omega=\{1,2,\ldots,n\}$ denote the vocabulary size under consideration. Note that the truncated linear program has two steps:
  • Sort the candidate tokens based on $q_{i}-p_{i}^{2}$. This requires $O(n\log n)$ computations. We select that $s$ largest tokens and identify this set as $\Omega_{1}$.
  • Apply linear program for the tokens in $\Omega_{1}$ which is the size of $s$. There are $O(s^{2})$ variables and standard implementation of the linear program involves a complexity of $O(s^{6})$.
Thus the overall complexity of our proposed scheme is $O(s^{6}+n\log n)$ where $n$ is the alphabet size. In practice we keep the size of $\Omega_{1}$ relatively small e.g., we set $s=5$ in many of our experiments. This avoids slowdown arising from the $s^{6}$ term in the linear program. Secondly, in practice, $\Omega$ is a subset of tokens from the original vocabulary after top-p sampling. So even when the original vocabulary is large, the value of $n$ in practice is relatively modest. In Appendix H we present the histogram of the alphabet size after top-p sampling to illustrate this point. Thus the $O(n\log n)$ is an acceptable cost for our implementation.
G.1 Variant of Truncated LP
We now present a variation of the LP that avoids the $O(n\log n)$ cost from sorting $\Omega$. The proposed method is as follows:
  • Select $s$ tokens with largest values of $q_{i}-p_{i}^{2}$. This requires $O(n\cdot s)$ computations. We identify this set as $\Omega_{1}$ and the put the remaining tokens in $\Omega_{2}$.
  • We propose the following heuristic choice of weights:
    $w_{i,j}=\begin{cases}1,&i\in\Omega_{1},j\in\Omega_{2}\\ 1/2,&i\in\Omega_{2},j\in\Omega_{2},i\neq j\end{cases}$ (195)
    Note that this results in the following distribution of the selected token:
    $p_{I}(k)=\begin{cases}p_{k}^{2}+\sum_{i=1,i\neq k}^{s}2p_{i}p_{k}w_{k,i}+\sum_{i=s+1}^{n}2p_{i}p_{k},&k\in\Omega_{1}\\ p_{k}^{2}+\sum_{i\neq k}^{n}p_{i}p_{k},&k\in\Omega_{2}\end{cases}$ (196)
    We leave the weights $w_{i,j}$ for $i<j$ and $i,j\in\Omega_{1}$ as free parameters.
  • Apply linear program for the tokens in $\Omega_{1}$ which is the size of $s$. There are $O(s^{2})$ variables and standard implementation of the linear program involves a complexity of $O(s^{6})$.
It can be seen that the overall complexity of the proposed method is $O(n\cdot s+s^{6})$. By following the same steps as in Section F can also be verified that the resulting probabilities in (196) are also satisfy (12) yielding the same theoretical guarantee.
Appendix H Histogram of Alphabet Size for OPT model
We present additional evidence that after top-p sampling the effective alphabet size can be significantly reduced. We consider the OPT model and report the histogram of the alphabet size under the draft and target models for the XSum task in Fig. 5, the Dolly task in Fig. 6, and the WMT task in Fig. 7. The histogram is truncated to $40$ tokens to make the figures clearer.
@@ Appendix H Histogram of Alphabet Size for OPT model
Appendix I LP and Fast LP version for non-identical draft distributions, $K=2$
@@
Appendix J Additional Experimental Results
J.1 Two Draft Models with Identical Temperatures
Here, we consider the case where identical draft models are used to generate candidate tokens sequences. We compare the performance of our method with SpecTr (Sun et al., 2024b) and SpecInfer (Miao et al., 2024), as well as single-draft speculative sampling (Leviathan et al., 2023, Chen et al., 2023), where we use the same temperature for the draft model.
In Figure 8, we set the temperature of the target model to $1.0$, while we vary the temperature of the $K=2$ draft models between the range of $1.2$ to $2.4$, and we report the performance achieved by the different schemes across the three tasks discussed earlier. The top plots report the block efficiency, which is the average number of tokens that are accepted per use of the draft model (Leviathan et al., 2023). The bottom plots show the percentage improvement in the token rate with respect to the baseline single draft scheme.
We observe that the IS scheme consistently outperforms the both SpecTr and SpecInfer across all three tasks. In fact when the temperature of the second draft is increased, the improvement in the block efficiency of both SpecTr and SpecInfer is rather negligible compared to the single draft baseline, while the token rate are in fact lower than the single draft baseline. On the other hand our proposed IS scheme is able to achieve consistent improvements in all the three tasks. In the experiments involving the XSum and WMT tasks we also measure the ROUGE-L (Lin, 2004) and BLEU (Papineni et al., 2002) scores respectively, which are reported in the Appendix J.3.
Scheme $K=2$ $K=3$ $K=4$ $K=5$ $K=6$
IS 2.13 $\pm$ 0.05 2.22 $\pm$ 0.05 2.26 $\pm$ 0.05 2.27 $\pm$ 0.05 2.28 $\pm$ 0.06
J.2 Three Draft Models
In this section, we consider the case of $K=3$ draft models that use different temperatures for token generation. As stated before, in case of non-identical draft models, the only plausible multi-draft sampling scheme for comparison is the SpecInfer scheme (Miao et al., 2024). We set the temperature of the target model to $1.0$, and the temperature of two of draft models to $1.2$ and $1.4$, while we vary the temperature of the first draft model between the range of $1.0$ to $1.6$. As observed in Figure 9, the proposed IS scheme outperforms SpecInfer and Single-draft speculative sampling method in terms of block efficiency and achieved token rate.
Figure 9: Performance comparison of different schemes on the Dolly dataset with $K=3$ drafts. We vary the temperature of the first draft model and keeping the temperature of the other two drafts to $1.2$ and $1.4$, with the target model having a temperature of $1.0$. The single-draft baseline uses a draft temperature of $1.2$.
J.3 ROUGE-L and BLEU Scores
@@
In the experiments involving the XSum and WMT tasks we measure the ROUGE-L (Lin, 2004) and BLEU (Papineni et al., 2002) scores respectively. Table 5 and Table 6 show the ROUGE-L and BLEU scores for the case of identical draft models, corresponding to the experiment in Figure 8 in Section J.1.
@@
Appendix K Notations
Table 9: Table of notations summarizing symbols and their descriptions.
Symbol Description
$x_{1:t}$ or $x_{1}^{t}$ Sequence of tokens $x_{1},x_{2},\cdots,x_{t}$
$K$ Number of draft models
$\Omega$ Vocabulary of tokens
$n$ Size of the vocabulary, $|\Omega|$
$\mathcal{S}$ The set of valid draft tokens under consideration at every step
$p$ Distribution of the draft model at any step, with $p_{i}$ denoting the probability of token $i$
$q$ Distribution of the target model at any step, with $q_{i}$ denoting the probability of token $i$
$\mathcal{M}_{s}$ Draft model
$\mathcal{M}_{b}$ Target model
$u^{t}$ Context sequence as a sequence of $t$ tokens in $\Omega^{t}$
$P^{*}(\text{acc})$ Optimal acceptance probability
$\beta_{y}(x_{1},\ldots,x_{K})$ Conditional probability that token $y$ is selected given the input tokens $x_{1},\ldots,x_{K}$
$\beta^{*}_{y}(x_{1},\ldots,x_{K})$ Optimal Conditional probability that token $y$ is selected given the input tokens $x_{1},\ldots,x_{K}$
$Y_{I}$ The selected token from the importance weighted sampling step
$Z$ The final output token after the combined importance weighted sampling
and speculative sampling step
$p_{I}(\cdot)$ Distribution of the output token in the importance weighted sampling step
$I$ Index of the selected token n the importance weighted sampling step
$w_{i,j}$ Conditional probability that token $i$ is selected given the input tokens are $i$ and $j$
$\alpha_{i,j}$ Joint probability that token $i$ is selected and the input tokens are $i$ and $j$
$\alpha_{i,j,k}$ Joint probability that token $i$ is selected and the input tokens are $i$, $j$, and $k$
$s$ LP-Truncation Threshold parameter
$\Omega_{0}$ High probability subset of tokens of the vocabulary $\Omega$